Operations Report – March 18th

Operations Report 18-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Tuesday Generator

8 am SOC was 70

5 pm SOC was 100

Generator was taken back to Grand Junction today and we are on solar power only.

Wednesday Generator

No longer using the generator. Solar Power only.

8 am SOC was 63

11 am SOC was 100

7 pm SOC was 97

Diesel Reading – 66%

Station Propane Reading – 69 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 10 gallons

Water Meter: n/a units

Water (static tank) – 200 gallons

Static to Loft Pump used – no

Water in GreenHab – basically empty gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: yes

Hours: 104.5

Beginning charge: 100%

Ending charge: 86%

Currently charging: yes

Spirit rover: no

Hours: 134.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: yes

Hours: 70.3

Beginning charge: 100%

Ending charge: 97%

Currently charging: yes

Curiosity rover used: yes

Hours: 139.6

Beginning charge: 100%

Ending charge: 96 %

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? Not used

CrewCar used and why, where? CrewCar is currently in Hanksville, used to deliver plants to Hanksville Community Gardens

General notes and comments: Removed more of the tunnel, hindered by strong winds, but wanted to get as many down as possible before rain. They are stored in the RAM until I can fold them.

Summary of internet: Nominal

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Only a few plants left in the herb garden. I am transplanting them into pots so they can be donated.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: It’s rainy tonight so we weren’t going to use the observatory but I asked Peter to make sure if the skies clear that it doesn’t not open. This way we will be able to determine how much of the nighttime use of power is me and how much is the observatory. Hopefully it will be okay once I am gone, but right now it looks like we use too much at night and will damage the batteries over the long term. We may need to stagger nights that we use the observatory so that every other day they don’t discharge so

much at night but that is a worse case scenario. I need to balance the daytime/nighttime power use so the station will not suffer any damage to the system once we are shut down.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: It’s raining again which means mud tomorrow. I’m so not happy about this.

Copyright © The Mars Society. All rights reserved. | Main Site